# A secure and private ensemble matcher using multi-vault obfuscated templates

### v1

@@ V Experimental setup
In our research, a face image is transformed into an embedding using five different deep learning methods (Fig. 4). Each of these embeddings is produced by a different convolutional neural network (CNN) model. In the current context, each of the embeddings ($e1,e2,e3,e4,e5$) is a 512 feature vector and numerically sufficiently different from each other. A total of five secure vaults is constructed. Each vault contains 2001 embeddings, of which 2000 embeddings are generated using 2000 random images of human faces and only 1 embedding is generated from an intended face image. In our research, we used a technique called StyleGAN2, developed by Karras et al [32] to generate realistic images of human faces. These images were then used to create random chaff points for the secure vault. [^1] To achieve this, we conducted an iterative process, generating 2000 embeddings from 2000 realistic face images obtained from the website thispersondoesnotexist.com This approach improves the privacy safeguards for the proposed method during the development and evaluation of the Multi-Vault Obfuscated Templates algorithm. The same structure, consisting of 2000 embeddings generated using random images and 1 embedding from an intended face image (i.e. the template), is replicated in all five vaults. To ensure the security, the hash value of the five different embeddings $H=hash(e1,e2,e3,e4,e5)$ is stored.
V-A Identification process
The process for a single vault is shown in Fig. 3. For a biometric identification match:
@@ V-A Identification process
  • Each query is compared against all embeddings across the five vaults, resulting in the calculation of a total of 2001*5 cosine similarity scores.
  • From each vault, the embedding with the highest cosine similarity score is selected, leading to the choice of five winners, one from each vault.
@@
  • Identification is matched only if this hash matches the previously generated hash (m) of the template.
@@
V-B Pre trained models
A set of five different pre-trained models have been used here to construct the end-to-end secure privacy framework. The first model is InceptionResnetV1; this particular model [33] has been pretrained on the VGGFace2dataset. [34, 35]. As a second classifier, another variation of InceptionResnetV1 has been used, but this version was pretrained with CASIA-Webface [36]. For both the models one and two, MTCNN [37] has been used as a preprocessing for face cropping. More precisely, MTCNN cropped a 250x250 dimension face to improve the model accuracy. As a third and fourth classifier, two different versions of insightFace [38] have been used. The third pre-trained classifier, named “buffalo”, is an implementation of the ResNet50 model pre-trained on the WebFace600K dataset. This dataset contains 600,000 unique identities and is derived from the larger WebFace42M dataset [39] which contains 42 million identities. The fourth pre-trained model antelopev2 is an implementation of ResNet100 pretrained using Glint360K [40], which contains 17091657 images of 360232 individuals. The fifth model is an arcface-based implementation of Resnet100 [41]; for this version some additional pre-processing has been used including image resizing, detection of faces using a face detector, face alignment, etc.
@@ VI Result and analysis
Our results are mainly described from two perspectives, accuracy and time complexity performance. Fig. 5 shows the comparison between the distribution of cosine similarity scores for genuine and imposter face images. The leftmost distribution is for the genuine cosine similarities, which mostly lie between 0.8 to 1. The middle distribution is for the imposter cosine similarity computed from the target template and other (non-mated) class images from a same dataset, with the score distribution mostly covering 0.2 to 0.4. The rightmost distribution also represents imposter cosine similarities but is computed by comparing the target templates and the random face images generated by GAN (as mentioned in the experimental setup), with the scores being mostly between -0.25 and 0.25. The constrained feature space of faces within a single dataset promotes higher cosine similarity (0.2-0.4) due to inherent structural similarities. Conversely, the inclusion of random external faces introduces greater vector variance, leading to a broader cosine similarity distribution (-0.25 to 0.25). These comparisons show how well the cosine similarity can distinguish the embeddings generated from face images. The corresponding ROC AUC of cosine similarity scores reinforces this fact.
Figure 5: Distribution of cosine similarity for Genuine, Other and Random classes
@@ VI-A Individual classifier performance
GT InceptionResnetV1-casia .9995
GT InsightFace Bufalo .9780
@@ VI-B End-to-end benchmarking
Figure 8: End-to-end performance based the on GT face dataset
Figure 9: End-to-end performance based the on LFW face dataset
@@ VI-C ROC Analysis
Figures 78, and 9 depict the study ROC curves for various datasets, all generated using the presented secure hash pipeline.Instead of evaluating individual classifiers, these ROC curves assess the performance of the complete registration pipeline that combines all five classifiers. Within this pipeline, a sample is considered a true positive if a genuine person is successfully registered, and a false positive otherwise. The ROC curves (Figures 789) depict the trade-off between security and accuracy in the registration process. These curves show how the rate of incorrectly accepting imposters (FPR) increases as we loosen the criteria for accepting genuine users (TPR). Imagine a vault containing 2001 images. The strictest system would only accept a single image with the highest cosine similarity score, while the most lenient would accept all 2001. Practically, accepting the top five highest-scoring images might be a good balance. This means the person can be authenticated if their true identity is among the top five closest matches out of the 2001 comparisons. As the ROC curves show, there’s a point beyond which accepting more images doesn’t significantly improve the chance of correctly identifying a genuine person, while increasing the risk of mistakenly accepting someone else.
@@
While benchmarking with various datasets, from Table III it can be seen that both the AT&T and GT face datasets yield an AUC greater than 0.99, whereas the LFW dataset yields an AUC greater than 0.9.
@@ VI-D Time complexity analysis
Image registration 1.15
Authentication 0.31
@@
Our system was developed and run on a Windows laptop, with an Intel i9 11th generation 64 bit CPU, 32 GB RAM, and 1 TB of storage. The system was equipped with an NVIDIA GeForce RTX 3080 GPU. As seen from Table IV, our method is computationally extremely efficient with an average of 1.15 seconds for image registration, 0.31 seconds for biometric matching, and 0.01 seconds of hash generation using the aforementioned system. Overall, the complete end-to-end procedure takes approximately 1.47 seconds on average.
@@ VII Conclusion
In this paper, we proposed a cryptographic biometric template security protocol similar to a fuzzy extractor. We showed that $2^{\gamma}$ security against brute-force attacks could be achieved by choosing appropriate parameters for the protocol. We tested our scheme against the AT&T, GT, and LFW datasets and showed that our protocol can preserve the average accuracy of the models used to generate the embeddings. We introduced a new approach to compute the receiver operating characteristic (ROC) curve which is compatible with our protocol. Our work leverages the power of generative models to create synthetic facial images, demonstrating their potential in building robust and secure end-to-end biometric system. Our approach offers distinct advantages as it eliminates the risk of exposing real individuals’ identities during the system’s training and operation.
Acknowledgments
This work was supported in part by grant from Center for Big Learning, an NSF IUCRC. Dr. Derakhshani is also a consultant for Jumio.

### v2

@@ V-A Experimental Setup
In our research, a face image is transformed into an embedding using five different deep learning methods (Fig. 3). Each of these embeddings is produced by a different convolutional neural network (CNN) model. The 512-dimensional feature vectors ($\boldsymbol{\mathit{e}}_{1},\boldsymbol{\mathit{e}}_{2},\boldsymbol{\mathit{e}}_{3},\boldsymbol{\mathit{e}}_{4},\boldsymbol{\mathit{e}}_{5}$) are numerically distinct, as verified by cosine similarity comparisons showing low similarity scores between vectors of different vaults. A total of five secure vaults is constructed. Each vault contains 2001 embeddings, of which 2000 embeddings are generated using 2000 random images of human faces and only 1 embedding is generated from an intended face image. We employed StyleGAN2 [24] to generate realistic human face images for our chaff points. StyleGAN2 was chosen for its superior image quality, diversity, and photo-realism compared to alternatives like PGGAN [23] and StarGAN [9]. Its improved fine-scale detail, reduced artifacts, and disentangled latent space make it ideal for creating indistinguishable and diverse chaff points for our secure vault system. These images were then used to create random chaff points for the secure vault. [^1] To achieve this, we conducted an iterative process, generating 2000 embeddings from 2000 realistic face images obtained from the website thispersondoesnotexist.com This approach improves the privacy safeguards for the proposed method during the development and evaluation of the Multi-Vault Obfuscated Templates algorithm. The same structure, consisting of 2000 embeddings generated using random images and 1 embedding from an intended face image (i.e. the template), is replicated in all five vaults. To ensure security, the hash value of the five different embeddings $H=hash(e1,e2,e3,e4,e5)$ is stored.
V-B Verification Process
The process for a single vault is shown in Fig. 2. For a biometric verification match:
@@ V-B Verification Process
  • Each resultant embedding is compared against all embeddings across the corresponding vault, resulting in the calculation of a total of 2001*5 cosine similarity scores.
  • The embedding with the highest cosine similarity score is selected from each vault, leading to the choice of five winners, one from each vault.
@@
  • The query is verified only if this hash matches the previously generated hash ( $H=hash(e1,e2,e3,e4,e5)$) of the template. However, in our implementation, we selected the top two or three embeddings from each vault and generated all possible $5^{2}$ or $5^{3}$ hashes, and compared them with the stored hash. This allows for relaxing the FRR of the system.
@@
V-C Pre-trained Models and Datasets
A set of five different pre-trained models have been used here to construct the end-to-end secure privacy framework. The first model is InceptionResnetV1; this particular model [41] has been pre-trained on the VGGFace2dataset. [5, 15]. As a second classifier, another variation of InceptionResnetV1 has been used, but this version was pre-trained with CASIA-Webface [50]. For both models, one and two, MTCNN [52] has been used as a preprocessing for face cropping. More precisely, MTCNN cropped a 250x250 dimension face to improve the model accuracy. As a third and fourth classifier, two different versions of insightFace [12] have been used. The third pretrained classifier, named “buffalo”, is an implementation of the ResNet50 model pretrained on the WebFace600K dataset. This dataset contains 600,000 unique identities and is derived from the larger WebFace42M dataset [54] which contains 42 million identities. The fourth pre-trained model antelopev2 is an implementation of ResNet100 pretrained using Glint360K [2], which contains 17091657 images of 360232 individuals. The fifth model is an arcface-based implementation of ResNet100 [18]; for this version, some additional preprocessing has been used including image resizing, detection of faces using a face detector, and face alignment.
@@ V-C Pre-trained Models and Datasets
Our method was evaluated using three widely recognized face datasets: the AT&T Database of Faces (AT&T), Labeled Faces in the Wild (LFW), and Georgia Tech Face Database (GT). AT&T contains 400 images of 40 subjects with controlled variations, LFW includes over 13,000 unconstrained face images collected from the web, and GT comprises 750 color images of 50 subjects with diverse facial expressions and illumination conditions. These datasets were selected to provide a comprehensive evaluation across different controlled and uncontrolled conditions, varying numbers of subjects, and diverse image characteristics, allowing us to assess our method’s performance in a range of scenarios.
@@ VI Result and Analysis
Our results are mainly described from two perspectives, accuracy and time complexity performance. Fig. 4 shows the comparison between the distribution of cosine similarity scores for genuine and imposter face images. The leftmost distribution is for the genuine cosine similarities, which mostly lie between 0.8 to 1. The middle distribution is for the imposter cosine similarity computed from the target template and other (non-mated) class images from the same dataset (e.g., same AT&T dataset but different identities), with the score distribution mostly covering 0.2 to 0.4. The rightmost distribution also represents imposter cosine similarities but is computed by comparing the target templates and the random face images generated by GAN (as mentioned in the experimental setup), with the scores being mostly between -0.25 and 0.25. The constrained feature space of faces within a single dataset promotes higher cosine similarity (0.2-0.4) due to inherent structural similarities. Conversely, the inclusion of random external faces introduces greater vector variance, leading to a broader cosine similarity distribution (-0.25 to 0.25). These comparisons show how well the cosine similarity can distinguish the embeddings generated from face images. The corresponding ROC AUC of cosine similarity scores reinforces this fact.
Figure 4: Distribution of Cosine Similarity for Genuine, Other, and Random Facial Images
@@ VI-A Individual Classifier Performance
GT InceptionReNnetV1-Casia .9995
GT InsightFace Buffalo .9780
@@ VI-B End-to-end Benchmarking
Figure 7: End-to-end Performance Based on the GT Face Dataset
Figure 8: End-to-end Performance Based on the LFW Face Dataset
@@ VI-C ROC Analysis
Figures 67, and 8 depict the study ROC curves for various datasets, all generated using the presented secure hash pipeline. Instead of evaluating individual classifiers, these ROC curves assess the performance of the complete biometric enrollment and authentication pipeline, which we refer to as the registration process. Within this pipeline, a sample is considered a true positive if a genuine person is successfully registered, and a false positive otherwise. The ROC curves (Figures 678) depict the trade-off between security and accuracy in the registration process. These curves show how the rate of incorrectly accepting imposters (FPR) increases as we loosen the criteria for accepting genuine users (TPR). Imagine a vault containing 2001 images. The strictest system would only accept a single image with the highest cosine similarity score, while the most lenient would accept all 2001. Practically, accepting the top five highest-scoring images might be a good balance. This means the person can be authenticated if their true identity is among the top five closest matches out of the 2001 comparisons. As the ROC curves show, there’s a point beyond which accepting more images doesn’t significantly improve the chance of correctly identifying a genuine person, while increasing the risk of mistakenly accepting someone else.
@@
As shown in Table III, the AT&T and Georgia Tech Face datasets achieve an AUC greater than 0.99, while the LFW dataset, with its challenging variability in pose, expression, and lighting, yields a lower AUC of 0.9042.
@@ VI-D Time Complexity Analysis
Face detection and embedding 1.15
Matching (cosine similarity) 0.31
@@
Our system, developed on a Windows laptop with an Intel i9 CPU and NVIDIA GeForce RTX 3080 GPU, demonstrates efficient performance. Face detection, preprocessing, and embedding creation take 1.15 seconds, matching 0.31 seconds, and secure hash generation 0.01 seconds(Table  IV). The process completes in an average of 1.47 seconds. Comparable non-protected systems report 0.1 to 0.5 seconds for feature extraction and matching. Our system maintains competitive performance while introducing additional security performance.
@@ VII Conclusion
In this paper, we proposed a cryptographic biometric template security protocol leveraging GenAI to enhance the security and privacy of face recognition. By harnessing GenAI’s ability to generate highly realistic and diverse synthetic data, we demonstrated a $2^{\gamma}$ security against brute-force attacks. Our experimental results on the AT&T, GT, and LFW datasets showed that our protocol preserves the average accuracy of the models used to generate the embeddings. We introduced a novel ROC curve computation method that evaluates our entire biometric pipeline, including multiple classifiers and GAN-generated chaff points. This comprehensive approach captures the synergistic effects of our multi-layered security system, unlike traditional ROC analyses that typically assess individual components separately. By utilizing GAN-generated synthetic data, we eliminated the risk of exposing real individuals’ identities during training and verification, hence providing additional privacy protections. This work showcases GenAI’s potential in constructing robust and secure end-to-end biometric systems. Future work may include exploring other GenAI methods to create more diverse samples and integrate them with other modalities, such as fingerprints, iris, and voice to develop multimodal biometric systems that are more resilient to attacks.
As GenAI continues its advance, one can envision a future where biometric security protocols continuously evolve to stay ahead of potential threats, enabling a new wave of secure and private biometric systems.
Acknowledgment
This work was supported in part by a grant from CBL, an NSF IUCRC. Dr. Derakhshani is also a consultant for Jumio.